# COREval: A Comprehensive and Objective Benchmark for Evaluating the Remote Sensing Capabilities of Large Vision-Language Models

### v1

@@ 7 Details on COREval Construction
For Image Caption, we randomly select 200 RSIs from Scene Classification ensuring an even distribution, and employ human annotators to generate a descriptive caption for each RSI. Then we prompt GPT-4 to generate 3 distracting descriptions, with these descriptions subsequently verified by human reviewers to ensure accuracy and relevance. The prompt we use is shown in Fig. 10.
For Landmark Recognition, we choose 100 most famous landmarks in the world from List Challenges [^6] https://www.listchallenges.com/ and download the RSIs with 0.3 meter/pixel spatial resolution from Google Earth according to their latitude and longitude coordinates.
For the other 6 leaf tasks in Single-instance Identification dimension, objection detection for identifying instances and obtaining their coordinates of bounding boxes is the first step to generate associated problems. As mentioned in Scene Classification, we also predefine a set of categories for instances based on the existing object detection datasets [44, 8] and OSM key-value pairs, as shown in Tab. 8. We select 30 of the 50 chosen cities as the primary data collection areas to capture more intra-class variations, including Riyadh, Gwangju, Tokyo, Daejeon, Chengdu, Istanbul, Nagoya, Wuhan, Hsinchu, Nanjing, Kuala Lumpur, Jeddah, Lyon, Paris, Rotterdam, Brussels, Berlin, Munich, Vienna, Zurich, Atlanta, Boston, Washington D.C., Montreal, Calgary, Windhoek, Gaborone, Rabat, Montevideo and Lima. Using the same procedure as for Scene Classification, we obtain the original RSIs from Google Earth and crop them into $800\times 800$ patches. We apply two object detection models, MTP [39] pretrained on DIOR-R [8] and YOLO [35] pretrained on DOTA [44], to automatically generate bounding boxes for the predefined categories. Human annotators are then employed to verify the completeness and accuracy of these labels, retaining only those with high confidence from both models. Once getting the labels and coordinates of these instances, we count the number of specific objects for Object Presence and Object Counting. We compute the centroids of these objects and assign their coarse locations in a $3\times 3$ grid for Object Localization, and use the normalized coordinates of bounding boxes directly as the label for Visual Grounding. For Hallucination Detection, we aim to assess the VLM’s capability to realize the non-existent instances within RSIs and correctly refuse to respond to such queries.
For Attribute Recognition, RSIs collected from Google Earth with a resolution of 0.3 meter/pixel are not sufficient for fine-grained attributes such as vehicle color. Therefore, we collect additional RSIs with a higher resolution of 0.1 meter/pixel. We then utilize K-means clustering to classify the color of the vehicles, generating MCQs related to this attribute.
For Change Detection, we collect registered RSIs captured at two different timestamps from 6 of the 50 cities, including Nicosia, Haifa, Dammam, Malmo, Seattle and Pretoria. Human annotators are enlisted to generate a descriptive caption for each pair of RSIs that highlights the temporal changes between them. GPT-4 is then prompted to generate 3 distracting options, with these descriptions subsequently verified by human reviewers to ensure accuracy and relevance. The prompt we use is shown in Fig. 10.
For Physical Property, we utilize the CNBH-10m [42] product, which provides the estimated building heights in China, to guide the collection of RSIs. We randomly select 100 points from this product and download the corresponding RSIs from Google Earth. To ensure accuracy, volunteers compare the estimated building heights provided by CNBH-10m with human-assessed heights from the RSIs. Any discrepancies are corrected by human annotators to maintain accuracy in building height measurements.
8 Details on Evaluation
By leveraging the instruction-following ability of VLMs, we restrict the model response to either the A/B/C/D option or the corresponding content of the choice. This constraint simplifies subsequent choice extraction and enables straightforward accuracy computation. As shown in Fig. 11, we design a system prompt tailored for both general-domain VLMs and RSVLMs when handling tasks with MCQs. For Visual Grounding, where the response requires free-form coordinates, we retain the original system prompts specific to RSVLMs, as these are better suited for generating precise coordinate outputs. This dual approach ensures optimal performance across varied task requirements while maintaining consistency and accuracy in evaluation.
We also present an example of evaluation strategy for CLIP-based VLMs in Fig. 11. The MCQ is transformed into $n$ declarative sentences by adding or replacing keywords in the question and further rearranging the word order. During the evaluation stage, the similarity score between the RSI and each sentence is computed. The choice corresponding to the sentence that achieves the highest similarity score is considered the model’s response. Notably, CLIP-based VLMs lack the capability for visual grounding, and directly computing similarity scores between RSIs and pure coordinates is not feasible. Consequently, this leaf task is left unaddressed for CLIP-based VLMs.
Failure Cases. For Visual Grounding, the responses from LHRS-Bot consistently fall within a limited set of repeated outputs across all problems, such as $(0.44,0.44,0.52,0.52)$, $(0.34,0.44,0.66,0.66)$ and $(0.38,0.38,0.59,0.59)$. This pattern results in extremely low accuracy and highlights a significant lack of diversity and adaptability in the model’s predictions for this task. For Attribute Comparison, all RSVLMs consistently output the choice “No”, regardless of how the choice order is altered. This behavior indicates a fundamental lack of capability in these models to perform attribute comparison tasks effectively.
9 Information of the Human Annotators
@@
We enlisted 12 human annotators with academic backgrounds in remote sensing or computer vision to deeply engage in the quality control stage of the COREval dataset. Their responsibilities included collecting suitable RSIs, verifying their attributes, constructing related problems, and ensuring the accuracy and relevance of all tasks and problems in COREval. Here is the specific information about these human annotators.
@@
Ethical Consideration. We explicitly informed the annotators about the intended use of the data and required them to ensure that the questions included in COREval were free from social bias, ethical issues, or privacy concerns throughout the annotation process.

### v2

@@ Appendix D Details on Evaluation
Appendix E More Experiment
E.1 Blind Evaluation
We claim to ensure the objectivity of our evaluation based on CHOICE by, on one hand, preventing data leakage, which is accomplished by collecting all data on our own from diverse platforms, and on the other hand, strictly ensuring that the correct answer can only be derived based on a thorough understanding of visual content. As pointed out in MMStar [13], visual content may be unnecessary for many evaluation samples, in which case the assessment of VLMs’ multi-modal capabilities degrades to merely evaluating the uni-modal performance of their LLM backbones. Therefore, we conducted a blind evaluation, where VLMs are required to answer the question without visual RSI input (considering the costs, we only conducted experiments on open-source models). Our expectation is that the results of VLMs across all tasks are similar to random selection (i.e., randomly choosing an answer as response).
Table 1011, and 12 present the outcomes of blind evaluation. To our expectation, the results for both L-2 and fine-grained L-3 tasks are close to the random selection, demonstrating the dependence on RSIs to answer the questions and confirming the reliability of our multi-modal benchmark. Notably, in the absence of visual input, some VLMs tend to respond with the refusal choice, such as “can’t judge” in Attribute Comparison (AC) and “The object is not in the picture” in Hallucination Detection (HD). Therefore, the high performances of the Qwen2-VL series, InternVL2 series and VHM in HD task of blind evaluation further demonstrate their strong HD capability, which alighs with the vanilla results in Table 4.
Table 10: Random and blind evaluation results for L-2 dimensions. Abbreviations adopted: ILC for Image-level Comprehension; SII for Single-instance Identification; CID for Cross-instance Discernment; AttR for Attribute Reasoning; AssR for Assessment Reasoning; CSR for common sense Reasoning.
Model ILC SII CID AttR AssR CSR
       Random 0.267 0.209 0.261 0.284 0.298 0.273
General-domain Large Vision-Language Models
Qwen2-VL-7B 0.253 0.299 0.292 0.238 0.153 0.246
Qwen2-VL-72B 0.280 0.304 0.178 0.142 0.215 0.290
Ovis1.6-Gemma2-9B 0.278 0.321 0.309 0.324 0.300 0.266
InternVL2-8B 0.262 0.355 0.286 0.280 0.290 0.320
InternVL2-26B 0.244 0.343 0.261 0.156 0.163 0.263
InternVL2-40B 0.270 0.283 0.318 0.244 0.075 0.296
LLaVA-1.6-7B 0.273 0.220 0.304 0.298 0.305 0.261
LLaVA-1.6-13B 0.268 0.258 0.341 0.332 0.268 0.277
Llama3.2-11B 0.219 0.234 0.157 0.214 0.173 0.293
GLM-4V-9B 0.275 0.257 0.328 0.312 0.345 0.223
DeepSeek-VL-7B 0.262 0.209 0.311 0.310 0.288 0.307
MiniCPM-V-2.5 0.284 0.293 0.279 0.322 0.338 0.223
Phi3-Vision 0.289 0.310 0.306 0.188 0.175 0.209
mPLUG-Owl3-7B 0.233 0.262 0.231 0.194 0.165 0.304
Molmo-7B-D 0.247 0.267 0.319 0.310 0.378 0.286
Remote Sensing Large Vision-Language Models
GeoChat 0.251 0.242 0.275 0.332 0.353 0.241
LHRS-Bot 0.276 0.172 0.165 0.320 0.375 0.243
LHRS-Bot-nova 0.258 0.270 0.287 0.198 0.133 0.291
VHM 0.255 0.390 0.289 0.248 0.100 0.240
Table 11: Fine-grained random and blind evaluation results for Perception. Abbreviations adopted: IM for Image Modality; IQ for Image Quality; MR for Map Recognition; SC for Scene Classification; IC for Image Caption; LR for Landmark Recognitionl; OC for Object Counting; OL for Object Localization; OP for Object Presence; AR for Attribute Recognition; HD for Hallucination Detection; AC for Attribute Comparison; SR for Spatial Relationship; CD for Change Detection.
  Image-Level Comprehension Single-Instance Identification Cross-Instance Relationship
Model IM IQ MR SC IC LR OC OL OP AR VG HD AC     SR CD
       Random 0.260 0.290 0.290 0.260 0.260 0.250 0.240 0.230 0.330 0.240 0.000 0.250 0.290     0.240 0.240
General-domain Large Vision-Language Models
Qwen2-VL-7B 0.250 0.170 0.250 0.280 0.320 0.320 0.070 0.360 0.080 0.500 0.000 0.840 0.320     0.240 0.320
Qwen2-VL-72B 0.240 0.180 0.270 0.330 0.350 0.240 0.060 0.410 0.070 0.360 0.000 1.000 0.000     0.290 0.320
Ovis1.6-Gemma2-9B 0.240 0.310 0.210 0.280 0.340 0.270 0.380 0.250 0.300 0.340 0.000 0.730 0.350     0.260 0.310
InternVL2-8B 0.260 0.310 0.250 0.240 0.300 0.250 0.480 0.340 0.120 0.490 0.000 0.790 0.310     0.270 0.270
InternVL2-26B 0.220 0.260 0.250 0.240 0.310 0.240 0.220 0.360 0.180 0.500 0.000 0.890 0.270     0.250 0.260
InternVL2-40B 0.270 0.230 0.250 0.280 0.340 0.250 0.290 0.290 0.190 0.550 0.000 0.440 0.360     0.280 0.300
LLaVA-1.6-7B 0.250 0.340 0.280 0.260 0.220 0.200 0.150 0.430 0.320 0.470 0.000 0.000 0.340     0.250 0.320
LLaVA-1.6-13B 0.240 0.320 0.280 0.260 0.250 0.240 0.160 0.470 0.340 0.550 0.010 0.070 0.370     0.300 0.350
Llama3.2-11B 0.240 0.170 0.270 0.220 0.290 0.190 0.360 0.330 0.110 0.390 0.000 0.270 0.050     0.200 0.280
GLM-4V-9B 0.280 0.300 0.220 0.270 0.240 0.230 0.240 0.360 0.310 0.450 0.000 0.240 0.350     0.300 0.330
DeepSeek-VL-7B 0.250 0.270 0.290 0.260 0.270 0.150 0.270 0.380 0.300 0.360 0.000 0.000 0.370     0.230 0.330
MiniCPM-V-2.5 0.270 0.330 0.260 0.270 0.310 0.240 0.240 0.360 0.230 0.460 0.000 0.540 0.310     0.290 0.210
Phi3-Vision 0.260 0.310 0.280 0.290 0.310 0.230 0.130 0.440 0.070 0.350 0.000 0.950 0.330     0.270 0.320
mPLUG-Owl3-7B 0.240 0.170 0.330 0.240 0.310 0.220 0.290 0.430 0.320 0.360 0.000 0.230 0.230     0.240 0.220
Molmo-7B-D 0.260 0.300 0.240 0.230 0.150 0.280 0.260 0.230 0.240 0.500 0.000 0.420 0.370     0.260 0.320
Remote Sensing Large Vision-Language Models
GeoChat 0.250 0.260 0.220 0.250 0.260 0.220 0.270 0.500 0.220 0.390 0.030 0.090 0.260     0.250 0.340
LHRS-Bot 0.240 0.290 0.250 0.290 0.250 0.200 0.210 0.120 0.200 0.330 - 0.020 -     0.260 0.310
LHRS-Bot-nova 0.210 0.230 0.270 0.280 0.330 0.190 0.320 0.240 0.320 0.460 0.000 0.350 0.340     0.210 0.310
VHM 0.250 0.230 0.290 0.260 0.290 0.220 0.350 0.370 0.250 0.500 0.050 0.920 0.300     0.270 0.300
Table 12: Fine-grained random and blind evaluation results for Reasoning. Abbreviations adopted: TP for Time Property; PP for Physical Property; EA for Environmental Assessment; RA for Resource Assessment; DD for Disaster Discrimination; GD for Geospatial Determination; SI for Situation Inference.
  Attribute Reasoning Assessment Reasoning Common Sense Reasoning
Model TP PP EA RA DD    GD SI
       Random 0.230 0.320 0.260 0.310 0.290    0.290 0.250
General-domain Large Vision-Language Models
Qwen2-VL-7B 0.250 0.230 0.220 0.130 0.240    0.260 0.240
Qwen2-VL-72B 0.220 0.090 0.470 0.130 0.260    0.260 0.330
Ovis1.6-Gemma2-9B 0.270 0.360 0.420 0.260 0.260    0.250 0.280
InternVL2-8B 0.310 0.260 0.140 0.340 0.310    0.240 0.380
InternVL2-26B 0.210 0.120 0.260 0.130 0.280    0.220 0.280
InternVL2-40B 0.220 0.260 0.150 0.050 0.240    0.300 0.330
LLaVA-1.6-7B 0.250 0.330 0.140 0.360 0.250    0.260 0.270
LLaVA-1.6-13B 0.230 0.400 0.170 0.300 0.230    0.260 0.320
Llama3.2-11B 0.220 0.210 0.390 0.100 0.250    0.250 0.350
GLM-4V-9B 0.270 0.340 0.210 0.390 0.220    0.200 0.240
DeepSeek-VL-7B 0.220 0.370 0.130 0.340 0.250    0.270 0.370
MiniCPM-V-2.5 0.250 0.370 0.330 0.340 0.260    0.280 0.160
Phi3-Vision 0.200 0.180 0.160 0.180 0.240    0.250 0.160
mPLUG-Owl3-7B 0.200 0.190 0.300 0.120 0.290    0.250 0.350
Molmo-7B-D 0.190 0.390 0.280 0.410 0.260    0.230 0.340
Remote Sensing Large Vision-Language Models
GeoChat 0.200 0.420 0.270 0.380 0.220    0.250 0.250
LHRS-Bot 0.260 0.360 0.240 0.420 0.290    0.200 0.240
LHRS-Bot-nova 0.180 0.210 0.350 0.060 0.300    0.240 0.320
VHM 0.260 0.240 0.280 0.040 0.270    0.180 0.260
E.2 Circular Evaluation
Figure 14: Circular Evaluation strategy. A problem is tested multiple times with circularly shifted choices and the VLM needs to succeed in all testing passes. In this example, the VLM failed in passes 2 and 4, thus considered failed the problem.
Following the configuration in MMBench [11], we also conducted a circular evaluation by circularly shifting the choices to give a more robust evaluation based on CHOICE. As shown in Figure 14, each question is fed to a VLM $N$ times ($N$ is the number of choices) and VLM is considered successful in answering the question only if it provides the correct answer in all circular passes.
Table 13: Circular evaluation results for L-2 dimensions. The negative values in the second column under each L-2 dimension represent the changes in performance after circular evaluation. The largest (second largest) change is in bold red (red). Abbreviations adopted: ILC for Image-level Comprehension; SII for Single-instance Identification; CID for Cross-instance Discernment; AttR for Attribute Reasoning; AssR for Assessment Reasoning; CSR for common sense Reasoning.
Model ILC SII CID AttR AssR CSR
       General-domain Large Vision-Language Models
[…]
Ovis1.6-Gemma2-9B 0.752 -0.077 0.543 -0.063 0.493 -0.139 -0.144 -0.144 -0.175 -0.175 0.841 -0.049
InternVL2-8B 0.701 -0.088 0.483 -0.083 0.490 -0.158 -0.164 -0.164 -0.225 -0.225 0.746 -0.070
[…]
LLaVA-1.6-7B 0.583 -0.103 0.367 -0.096 0.349 -0.224 -0.190 -0.190 -0.215 -0.215 0.634 -0.114
LLaVA-1.6-13B 0.639 -0.081 0.415 -0.107 0.498 -0.128 -0.172 -0.172 -0.123 -0.123 0.661 -0.071
Llama3.2-11B 0.608 -0.137 0.341 -0.161 0.224 -0.280 -0.182 -0.182 -0.373 -0.373 0.686 -0.117
GLM-4V-9B 0.649 -0.135 0.448 -0.107 0.486 -0.158 -0.212 -0.212 -0.300 -0.300 0.783 -0.080
DeepSeek-VL-7B 0.687 -0.077 0.488 -0.073 0.428 -0.166 -0.088 -0.088 -0.075 -0.075 0.801 -0.039
MiniCPM-V-2.5 0.633 -0.124 0.428 -0.119 0.424 -0.207 -0.136 -0.136 -0.175 -0.175 0.770 -0.087
Phi3-Vision 0.603 -0.107 0.398 -0.104 0.409 -0.172 -0.160 -0.160 -0.143 -0.143 0.617 -0.083
mPLUG-Owl3-7B 0.656 -0.109 0.412 -0.112 0.226 -0.261 -0.214 -0.214 -0.230 -0.230 0.786 -0.064
Molmo-7B-D 0.590 -0.131 0.356 -0.197 0.384 -0.264 -0.172 -0.172 -0.205 -0.205 0.613 -0.111
Remote Sensing Large Vision-Language Models
GeoChat 0.501 -0.142 0.365 -0.104 0.251 -0.229 -0.202 -0.202 -0.274 -0.274 0.570 -0.127
LHRS-Bot 0.462 -0.171 0.150 -0.141 0.005 -0.166 -0.222 -0.222 -0.426 -0.426 0.393 -0.217
LHRS-Bot-nova 0.478 -0.210 0.261 -0.268 0.240 -0.286 -0.336 -0.336 -0.120 -0.120 0.341 -0.303
VHM 0.634 -0.117 0.589 -0.114 0.180 -0.256 -0.266 -0.266 -0.256 -0.256 0.541 -0.203
Table 14: Fine-grained circular evaluation results for Perception. Abbreviations adopted: IM for Image Modality; IQ for Image Quality; MR for Map Recognition; SC for Scene Classification; IC for Image Caption; LR for Landmark Recognitionl; OC for Object Counting; OL for Object Localization; OP for Object Presence; AR for Attribute Recognition; HD for Hallucination Detection; AC for Attribute Comparison; SR for Spatial Relationship; CD for Change Detection.
  Image-Level Comprehension Single-Instance Identification Cross-Instance Relationship
Model IM IQ MR SC IC LR OC OL OP AR HD AC     SR CD
       General-domain Large Vision-Language Models
[…]
InternVL2-8B 0.590 0.370 0.250 0.900 0.830 0.710 0.300 0.440 0.850 0.440 0.710 0.750     0.180 0.500
[…]
LLaVA-1.6-13B 0.420 0.340 0.180 0.870 0.760 0.620 0.150 0.530 0.790 0.480 0.220 0.770     0.260 0.380
Llama3.2-11B 0.530 0.090 0.480 0.840 0.780 0.790 0.270 0.330 0.820 0.420 0.180 0.260     0.090 0.360
[…]
MiniCPM-V-2.5 0.450 0.160 0.500 0.880 0.900 0.750 0.290 0.360 0.880 0.480 0.510 0.650     0.130 0.470
Phi3-Vision 0.390 0.270 0.250 0.840 0.700 0.520 0.110 0.670 0.870 0.400 0.270 0.660     0.250 0.210
mPLUG-Owl3-7B 0.600 0.220 0.480 0.850 0.830 0.880 0.320 0.520 0.910 0.370 0.250 0.210     0.110 0.430
Molmo-7B-D 0.420 0.350 0.140 0.770 0.780 0.540 0.160 0.400 0.820 0.450 0.320 0.650     0.090 0.360
Remote Sensing Large Vision-Language Models
GeoChat 0.118 0.068 0.255 0.857 0.395 0.710 0.012 0.644 0.928 0.256 0.000 0.463     0.073 0.145
LHRS-Bot 0.010 0.043 0.006 0.841 0.520 0.200 0.000 0.000 0.888 0.030 $0$     0.000 0.020
LHRS-Bot-nova 0.084 0.021 0.006 0.849 0.550 0.420 0.000 0.118 0.926 0.154 0.066 0.446     0.000 0.240
[…]
Table 15: Fine-grained circular evaluation results for Reasoning. Abbreviations adopted: TP for Time Property; PP for Physical Property; EA for Environmental Assessment; RA for Resource Assessment; DD for Disaster Discrimination; GD for Geospatial Determination; SI for Situation Inference.
  Attribute Reasoning Assessment Reasoning Common Sense Reasoning
Model TP PP EA RA DD    GD SI
       General-domain Large Vision-Language Models
Qwen2-VL-7B 0.420 0.560 0.130 0.370 0.850    0.950 0.770
[…]
InternVL2-8B 0.290 0.640 0.010 0.470 0.840    0.720 0.700
[…]
LLaVA-1.6-7B 0.020 0.420 0.090 0.260 0.660    0.570 0.660
LLaVA-1.6-13B 0.110 0.430 0.310 0.360 0.700    0.550 0.710
Llama3.2-11B 0.170 0.350 0.030 0.010 0.720    0.810 0.580
GLM-4V-9B 0.100 0.520 0.030 0.190 0.810    0.940 0.660
DeepSeek-VL-7B 0.290 0.540 0.110 0.360 0.840    0.840 0.750
[…]
Phi3-Vision 0.190 0.320 0.110 0.110 0.800    0.430 0.620
mPLUG-Owl3-7B 0.070 0.300 0.260 0.060 0.880    0.760 0.740
[…]
Remote Sensing Large Vision-Language Models
GeoChat 0.035 0.280 0.000 0.125 0.555    0.720 0.480
LHRS-Bot 0.000 0.240 0.000 0.000 0.390    0.135 0.567
LHRS-Bot-nova 0.020 0.177 0.000 0.000 0.280    0.165 0.500
VHM 0.195 0.080 0.080 0.095 0.650    0.515 0.487
Table 1314 and 15 shows the results of circular evaluation. The performance of each model has declined to varying degrees, with RSVLMs showing the most significant drop. Specifically, LHRS-Bot-nova experiences the largest drop in 4 out of 6 L-2 dimensions (ILC, CIR, AttR and CSR), while LHRS-Bot shows a similar decline in 3 out of 6 L-2 dimensions (ILC, AssR and CSR). An in-depth analysis of L-3 tasks reveals that several tasks have an accuracy near or even equals to 0, indicating the weak capability in specific remote sensing tasks, such as OC, HD, SR, EA and RA. This phenomenon suggests that RSVLMs are heavily influenced by different prompts, leading to reduced feasibility. Qwen2-VL-72B and InternVL2-40B continue to secure the top two places as shown in the vanilla results and have a relatively smaller performance drop, further highlighting their strong and robust capabilities in remote sensing tasks. From the perspective of L-2 dimensions, both AttR and AssR have the greatest negative impact on the performance of each VLM, indicating that when faced with problems requiring reasoning based on the perceived visual content, VLMs are significantly less capable and exhibit a high degree of randomness. Therefore, there is an urgent need for further improvement in reasoning capabilities of VLMs.
Appendix F Quality Control
F.1 Quality Control Procedure
We implemented a multi-stage quality control process by recruiting 12 volunteers with professional backgrounds in remote sensing or computer vision, including both master’s and doctoral students, to assist in the construction and quality assurance of CHOICE. These volunteers were randomly divided into two groups of six.
During data collection and question creation, two of the six volunteers gather the required RSIs from various data sources, while another two extract and verify the attributes and detailed information of these RSIs. The remaining two take the lead in preliminarily constructing questions based on this information, ensuring that each question aligns accurately with the targeted capabilities.
During the quality inspection process, the six volunteers were further equally divided into two teams, with each team responsible for verifying the quality of specific tasks within CHOICE. A question is approved only if all three team members confirmed it was error-free. If one or more volunteers identified an issue, it would be returned to the first group for further discussion and revision.
F.2 Information of the Human Annotators
@@
We enlisted 12 human annotators with academic backgrounds in remote sensing or computer vision to deeply engage in the quality control stage of the CHOICE dataset. Their responsibilities included collecting suitable RSIs, verifying their attributes, constructing related problems, and ensuring the accuracy and relevance of all tasks and problems in CHOICE. Here is the specific information about these human annotators.
@@
Ethical Consideration. We explicitly informed the annotators about the intended use of the data and required them to ensure that the questions included in CHOICE were free from social bias, ethical issues, or privacy concerns throughout the annotation process.